please help me in finding the Internal angles of a Triangle when the external angle is given as 75
Answer by rothauserc(4718)

we have two equations, let x, y, z be the internal angles of the triangle and w be the external angle with z, then we have
x + y + z = 180
w + z = 180
solve second equation for z
z = 180 - w
now substitute for z in first equation
x + y + (180 - w) = 180
x + y - w = 0
w = x + y
we know w = 75 degrees and
z + 75 = 180
z = 105 and x+y = 75